# Linear and nonlinear instability of vortex columns

### v2

@@ 3.4.1. Constructing approximate solutions
$\omega(\cdot,t)=\omega_{0}-\int_{-\infty}^{t}\mathrm{e}^{(t-s)L}[u,\omega]\,\mathrm{d}s\,,$ (3.45)
with $\omega-\omega_{0}=O(e^{2ta})$ in a certain topology. The integral equations lose one derivative, but this will not be an issue, as the eigenfunction is smooth, and we will consider finitely many $k$’s.
To be precise we first note that the assumptions of Theorem 1.2 imply not only that $L$ has an unstable eigenvalue but also that one of the invariant subspaces must contain an unstable mode. Let $\lambda\in\sigma(L_{n,\alpha})$ be an unstable eigenvalue of $L_{n,\alpha}$ for some $(n,\alpha)\in\mathbb{Z}^{2}\setminus\{(\pm 1,0)\}$ satisfying
@@
$\omega_{0}(\cdot,t):=\mathrm{e}^{tL}\eta+\overline{\mathrm{e}^{tL}\eta}=\mathrm{e}^{\lambda t}\eta+\mathrm{e}^{\overline{\lambda}t}\overline{\eta},$ (3.47)
@@
$\partial_{t}\omega_{0}-L\omega_{0}=0.$ (3.48)
@@
$\omega_{k}(\cdot,t):=\omega_{0}-\int_{-\infty}^{t}\mathrm{e}^{(t-s)L}[u_{k-1},\omega_{k-1}]\,\mathrm{d}s.$ (3.50)
@@
$\left\|\int_{-\infty}^{t}\mathrm{e}^{(t-s)L}[u_{k-1},\omega_{k-1}]\mathrm{d}s\right\|_{L^{Q}\cap H^{m}}$ $\leq C\int_{-\infty}^{t}\mathrm{e}^{S(t-s)}\|[u_{k-1},\omega_{k-1}](\cdot,s)\|_{L^{Q}\cap H^{m}}\,\mathrm{d}s$ (3.52)
@@
where $Q^{*}_{2}:=2Q/(2-Q)$ denotes the $2$D Sobolev exponent of $Q$. This in particular shows (3.43).
@@ 3.4.2. Nonlinear estimate
$\partial_{t}{\widetilde{\omega}}-L{\widetilde{\omega}}+[{\widetilde{u}},{\widetilde{\omega}}]+[{\widetilde{u}},{\omega_{k}}]+[{u_{k}},{\widetilde{\omega}}]+\underbrace{[u_{k}-u_{k-1},\omega_{k}]+[u_{k-1},\omega_{k}-\omega_{k-1}]}_{=:-E_{k}}=0$ (3.55)
@@
satisfies the $3$D Euler equations (1.1). We note that, by classical theory for the solution $\omega={\overline{\omega}}+\omega_{k}+{\widetilde{\omega}}^{(\ell)}$, a solution exists on a short time interval which a priori may depend on $\ell$. Our goal in this step is to estimate ${\widetilde{\omega}}^{(\ell)}$ and thereby demonstrate that it exists up to time $T=O(1)$ and decays like $O(\mathrm{e}^{a(k+1)t})$ backward in time.
@@
[…]
@@ 3.4.3. Conclusion
where $\omega_{0}=\mathrm{e}^{\lambda t}\eta+\mathrm{e}^{\overline{\lambda}t}\overline{\eta}$ is explicit and satisfies
@@
$\begin{split}\|u(\overline{T})-{\overline{u}}\|_{L^{\infty}}&\geq\|u_{0}(\overline{T})\|_{L^{\infty}}-\|(u_{k}-u_{0})(\overline{T})\|_{L^{\infty}}-\|{\widetilde{u}}(\overline{T})\|_{L^{\infty}}\geq\widetilde{C}^{-1}\mathrm{e}^{a\overline{T}}-\widetilde{C}\mathrm{e}^{2a\overline{T}}-\widetilde{C}\mathrm{e}^{2a\overline{T}}\\ &\geq\frac{1}{2\widetilde{C}}\mathrm{e}^{a\overline{T}}=:\delta\,,\end{split}$ (3.74)
@@ Appendix A Expansion of $k$
$\begin{split}k(r_{0})&=p_{0}n^{2}\left(1+\frac{b_{0}}{(-ib_{0}^{1/2}-{\overline{\omega}})^{2}}+\frac{a_{0}}{n(-ib_{0}^{1/2}-{\overline{\omega}})}+\frac{d_{0}}{n^{2}}\right)\\ &=p_{0}n^{2}\left(-i{\overline{\omega}}\frac{2}{b_{0}^{1/2}}+{{\overline{\omega}}}^{2}\frac{3}{b_{0}}+O({\overline{\omega}}^{3})+i\frac{a_{0}}{nb_{0}^{1/2}}+O({\overline{\omega}}n^{-1})+O(n^{-2})\right),\\ k^{\prime}(r_{0})&=p^{\prime}(r_{0})n^{2}\left(-i{\overline{\omega}}\frac{2}{b_{0}^{1/2}}+{{\overline{\omega}}}^{2}\frac{3}{b_{0}^{2}}+O({\overline{\omega}}^{3})+O(n^{-1})\right)+p_{0}n^{2}\left(\frac{a^{\prime}(r_{0})}{n(-ib_{0}^{1/2}-{\overline{\omega}})}+\frac{d^{\prime}(r_{0})}{n^{2}}\right)\\ &=n^{2}\left(-i{\overline{\omega}}\frac{2p^{\prime}(r_{0})}{b_{0}^{1/2}}+O({\overline{\omega}}^{2})+O(n^{-1})\right)\\ k^{\prime\prime}(r_{0})&=p^{\prime\prime}(r_{0})n^{2}\left(-i{\overline{\omega}}\frac{2}{b_{0}^{1/2}}+{{\overline{\omega}}}^{2}\frac{3}{b_{0}^{2}}+O({\overline{\omega}}^{3})+O(n^{-1})\right)+2p^{\prime}(r_{0})n^{2}\left(\frac{a^{\prime}(r_{0})}{n(-ib_{0}^{1/2}-{\overline{\omega}})}+\frac{d^{\prime}(r_{0})}{n^{2}}\right)\\ &\quad+p_{0}n^{2}\left(\frac{b^{\prime\prime}(r_{0})}{(-ib_{0}^{1/2}-{\overline{\omega}})^{2}}-\frac{2b_{0}n\Lambda^{\prime\prime}(r_{0})}{(-ib_{0}^{1/2}-{\overline{\omega}})^{3}}+\frac{a^{\prime\prime}(r_{0})}{n(-ib_{0}^{1/2}-{\overline{\omega}})}-\frac{a(r_{0})n\Lambda^{\prime\prime}(r_{0})}{n(-ib_{0}^{1/2}-{\overline{\omega}})^{2}}+\frac{d^{\prime\prime}(r_{0})}{n^{2}}\right)\\ &=n^{2}\left(-i{\overline{\omega}}\frac{2p^{\prime\prime}(r_{0})}{b_{0}^{1/2}}+O({\overline{\omega}}^{2})+O(n^{-1})-\frac{p_{0}b^{\prime\prime}(r_{0})}{b_{0}}-i{\overline{\omega}}\frac{2p_{0}b^{\prime\prime}(r_{0})}{b_{0}^{3/2}}+in\frac{2p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}^{1/2}}\right.\\ &\hskip 113.81102pt\left.+n{\overline{\omega}}\frac{6p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+\frac{a(r_{0})p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+i{\overline{\omega}}\frac{2p_{0}a(r_{0})\Lambda^{\prime\prime}(r_{0})}{b_{0}^{3/2}}\right)\end{split}$ (A.1)
Note that the leading order expansion of $k^{\prime}(r_{0})$ (arising from taking $\mathrm{d}/\mathrm{d}r$ of $b/\gamma^{2}$) vanishes due to our choice (1.14) of $r_{0},\beta$.
@@
$|\widetilde{\omega}|+|r-r_{0}|\leq Bn^{-1/2}$ (A.2)
(recall (1.27) concerning $\widetilde{\omega}$) for $n$ sufficiently large, then
@@
where the implicit constant in the $O(1)$ may depend on $B$. (Recall (2.6) that $k_{\rm err}(r)=k(r)-(k_{0}+k_{2}(r-r_{0})^{2})$.) To this end, we take into account the assumption ${\widetilde{\omega}}=O(n^{-1/2})$ from (A.2) to further reduce (A.1) to
$\begin{split}k(r_{0})&=k_{0}-in^{2}{\widehat{\omega}}\frac{2p_{0}}{b_{0}^{1/2}}+n^{2}{\widetilde{\omega}}^{2}\frac{3p_{0}}{b_{0}}+\widehat{\omega}O(n^{3/2})+in\frac{a_{0}}{b_{0}^{1/2}}+O(n^{1/2})=k_{0}+O(n^{3/2}),\\ k^{\prime}(r_{0})&=(-1+i)O(n^{3/2}),\\ k^{\prime\prime}(r_{0})/2&=k_{2}+n^{3}{\widetilde{\omega}}\frac{3p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+n^{3}{\widehat{\omega}}\frac{3p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+O(n^{2}),\end{split}$ (A.4)
@@
We verify that, in the case of the trailing vortex (1.3), if $q$ is sufficiently small, then
@@
Note that this equation has no solution $r>0$ if $\beta\leq q/2$ (in that case (A.8), with “$=$” replaced by $\geq$”, holds for all $r\geq 0$, with equality only for $r=0$). On the other hand for each $\beta>q/2$ (A.8) has a unique $r_{0}>0$. Clearly the solution $r_{0}=r_{0}(\beta)$ is an increasing bijection $r_{0}\colon(q/2,\infty)\to(0,\infty)$.
On the other hand, for the trailing vortex (1.3) we have $b(r)=4\beta(1-\beta q)q(1-\mathrm{e}^{-r^{2}})\mathrm{e}^{-r^{2}}/(1+\beta^{2}r^{2})$, and $b^{\prime}(r)=0$ if and only if
$g(r)-\beta^{2}=0,\qquad\text{ where }g(r):=\frac{1-2\mathrm{e}^{-r^{2}}}{r^{2}(2\mathrm{e}^{-r^{2}}-1)+\mathrm{e}^{-r^{2}}-1}.$ (A.9)
In particular $b$ has a sign, depending on the sign of $1-\beta q$, and so the third condition in (1.14) implies that we must have $\beta<1/q$. This together with the lower bound on $\beta$ above implies the constraint (1.15).
This in particular gives the first smallness requirement on $q$, $q<2^{1/2}$.
We now observe that $g(r)>0$ if and only if $r<(\log 2)^{1/2}$, and that $g$ is decreasing for such $r$ with $g(r)\to+\infty$ as $r\to 0^{+}$. We thus let $q$ be sufficiently small so that ($q<2^{1/2}$ and) the solution $r$ to (A.8) with $\beta=1/q$ satisfies $r>\log 2$, and we let $\beta_{0}\in(q/2,1/q)$ be such that $\log 2$ is the solution to (A.8) with $\beta=\beta_{0}$. This implies that $g\circ r_{0}\colon(q/2,\beta_{0})\to(0,\infty)$ is a decreasing bijection, and so $g\circ r_{0}-(\cdot)^{2}\colon(q/2,\beta_{0})\to(-\beta_{0}^{2},\infty)$ is a decreasing bijection. Hence there exists a unique $\beta\in(q/2,\beta_{0})$ such that $g(r_{0}(\beta))-\beta^{2}=0$. Moreover, $\beta$ is unique also on $(q/2,1/q)$, as $g(r_{0}(\beta))\leq 0$ for $\beta\geq\beta_{0}$.
It is also clear that, for such $r_{0},\beta$, we have that $b(r_{0})$ is the global maximum of $b$ and $\Lambda(r_{0})$ is the global minimum of $\Lambda$. Thus we obtain (A.7), as required.
As for the upper bound in (1.15), we also note that $b=0$ if $\beta=1/q$, in which case $\mathrm{Re}\,k>0$ for all $r$ and there are no nontrivial solutions to (1.17).

### v3

@@ 3.5.1. Constructing approximate solutions
$\omega(\cdot,t)=\omega_{0}-\int_{-\infty}^{t}\mathrm{e}^{(t-s)\mathcal{L}}[u,\omega]\,\mathrm{d}s\,$, (3.45)
with $\omega-\omega_{0}=O(\mathrm{e}^{2ta})$ in a certain topology. The integral equations lose one derivative, but this will not be an issue, as the eigenfunction is smooth, and we will consider finitely many $k$’s.
To be precise we first note that the assumptions of Theorem 1.2 imply not only that $\mathcal{L}$ has an unstable eigenvalue but also that one of the invariant subspaces must contain an unstable mode. Let $\lambda\in\sigma(L_{n,\alpha})$ be an unstable eigenvalue of $L_{n,\alpha}$ for some $(n,\alpha)\in\mathbb{Z}^{2}\setminus\{(\pm 1,0)\}$ satisfying
@@
$\omega_{0}(\cdot,t):=\mathrm{e}^{t\mathcal{L}}\eta+(\mathrm{e}^{t\mathcal{L}}\eta)^{*}=\mathrm{e}^{\lambda t}\eta+\mathrm{e}^{\lambda^{*}t}\eta^{*},$ (3.47)
@@
$\partial_{t}\omega_{0}-\mathcal{L}\omega_{0}=0.$ (3.48)
@@
$\omega_{k}(\cdot,t):=\omega_{0}-\int_{-\infty}^{t}\mathrm{e}^{(t-s)\mathcal{L}}[u_{k-1},\omega_{k-1}]\,\mathrm{d}s.$ (3.50)
@@
$\left\|\int_{-\infty}^{t}\mathrm{e}^{(t-s)\mathcal{L}}[u_{k-1},\omega_{k-1}]\mathrm{d}s\right\|_{L^{Q}\cap H^{m}}$ $\leq C\int_{-\infty}^{t}\mathrm{e}^{S(t-s)}\|[u_{k-1},\omega_{k-1}](\cdot,s)\|_{L^{Q}\cap H^{m}}\,\mathrm{d}s$ (3.52)
@@
where $Q^{*}_{2}:=2Q/(2-Q)$ denotes the 22D Sobolev exponent of $Q$. This in particular shows (3.43).
@@ 3.5.2. Nonlinear estimate
$\partial_{t}{\widetilde{\omega}}-\mathcal{L}{\widetilde{\omega}}+[{\widetilde{u}},{\widetilde{\omega}}]+[{\widetilde{u}},{\omega_{k}}]+[{u_{k}},{\widetilde{\omega}}]+\underbrace{[u_{k}-u_{k-1},\omega_{k}]+[u_{k-1},\omega_{k}-\omega_{k-1}]}_{=:-E_{k}}=0$ (3.55)
@@
satisfies the 33D Euler equations (1.1). We note that, by classical theory for the solution $\omega={\overline{\omega}}+\omega_{k}+{\widetilde{\omega}}^{(\ell)}$, a solution exists on a short time interval which a priori may depend on $\ell$. Our goal in this step is to estimate ${\widetilde{\omega}}^{(\ell)}$ and thereby demonstrate that it exists up to time $T=O(1)$ and decays like $O(\mathrm{e}^{a(k+1)t})$ backward in time.
@@
$\int f\cdot|{\widetilde{\omega}}|^{Q-2}{\widetilde{\omega}}\,dx\leq\|f\|_{L^{Q}}\|{\widetilde{\omega}}^{Q-1}\|_{L^{Q/(Q-1)}}=\|f\|_{L^{Q}}\|{\widetilde{\omega}}\|_{L^{Q}}^{Q-1}\leq C_{\varepsilon}\|f\|_{L^{Q}}^{Q}+\varepsilon\|{\widetilde{\omega}}\|_{L^{Q}}^{Q}\,$, (3.63)
@@ 3.5.3. Conclusion
where $\omega_{0}=\mathrm{e}^{\lambda t}\eta+\mathrm{e}^{\lambda^{*}t}\eta^{*}$ is explicit and satisfies
@@
[…]
@@ Appendix A Expansion of $k$
$\begin{split}k(r_{0})&=p_{0}n^{2}\left(1+\frac{b_{0}}{(-ib_{0}^{1/2}-\mu_{m})^{2}}+\frac{a_{0}}{n(-ib_{0}^{1/2}-\mu_{m})}+\frac{d_{0}}{n^{2}}\right)\\ &=p_{0}n^{2}\left(-i\mu_{m}\frac{2}{b_{0}^{1/2}}+\mu_{m}^{2}\frac{3}{b_{0}}+O(\mu_{m}^{3})+i\frac{a_{0}}{nb_{0}^{1/2}}+O(\mu_{m}n^{-1})+O(n^{-2})\right),\\ k^{\prime}(r_{0})&=p^{\prime}(r_{0})n^{2}\left(-i\mu_{m}\frac{2}{b_{0}^{1/2}}+\mu_{m}^{2}\frac{3}{b_{0}^{2}}+O(\mu_{m}^{3})+O(n^{-1})\right)+p_{0}n^{2}\left(\frac{a^{\prime}(r_{0})}{n(-ib_{0}^{1/2}-\mu_{m})}+\frac{d^{\prime}(r_{0})}{n^{2}}\right)\\ &=n^{2}\left(-i\mu_{m}\frac{2p^{\prime}(r_{0})}{b_{0}^{1/2}}+O(\mu_{m}^{2})+O(n^{-1})\right)\\ k^{\prime\prime}(r_{0})&=p^{\prime\prime}(r_{0})n^{2}\left(-i\mu_{m}\frac{2}{b_{0}^{1/2}}+\mu_{m}^{2}\frac{3}{b_{0}^{2}}+O(\mu_{m}^{3})+O(n^{-1})\right)+2p^{\prime}(r_{0})n^{2}\left(\frac{a^{\prime}(r_{0})}{n(-ib_{0}^{1/2}-\mu_{m})}+\frac{d^{\prime}(r_{0})}{n^{2}}\right)\\ &\quad+p_{0}n^{2}\left(\frac{b^{\prime\prime}(r_{0})}{(-ib_{0}^{1/2}-\mu_{m})^{2}}-\frac{2b_{0}n\Lambda^{\prime\prime}(r_{0})}{(-ib_{0}^{1/2}-\mu_{m})^{3}}+\frac{a^{\prime\prime}(r_{0})}{n(-ib_{0}^{1/2}-\mu_{m})}-\frac{a(r_{0})n\Lambda^{\prime\prime}(r_{0})}{n(-ib_{0}^{1/2}-\mu_{m})^{2}}+\frac{d^{\prime\prime}(r_{0})}{n^{2}}\right)\\ &=n^{2}\left(-i\mu_{m}\frac{2p^{\prime\prime}(r_{0})}{b_{0}^{1/2}}+O(\mu_{m}^{2})+O(n^{-1})-\frac{p_{0}b^{\prime\prime}(r_{0})}{b_{0}}-i\mu_{m}\frac{2p_{0}b^{\prime\prime}(r_{0})}{b_{0}^{3/2}}+in\frac{2p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}^{1/2}}\right.\\ &\hskip 113.81102pt\left.+n\mu_{m}\frac{6p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+\frac{a(r_{0})p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+i\mu_{m}\frac{2p_{0}a(r_{0})\Lambda^{\prime\prime}(r_{0})}{b_{0}^{3/2}}\right)\end{split}$ (A.1)
Note that the leading order terms of $k^{\prime}(r_{0})$ (arising from taking $\mathrm{d}/\mathrm{d}r$ of $b/\gamma^{2}$) vanish due to our choice (1.14) of $r_{0},\beta$.
@@
$|\mu_{m}|+|r-r_{0}|\leq Bn^{-1/2}$ (A.2)
(recall (1.28) concerning $\mu_{m}$) for $n$ sufficiently large, then
@@
where the implicit constant in the $O(1)$ may depend on $B$. (Recall (2.6) that $k_{\rm err}(r)=k(r)-(k_{0}+k_{2}(r-r_{0})^{2})$.) To this end, we take into account the assumption $\mu_{m}=O(n^{-1/2})$ from (A.2) to further reduce (A.1) to
$\begin{split}k(r_{0})&=k_{0}-in^{2}{\widehat{\omega}}\frac{2p_{0}}{b_{0}^{1/2}}+n^{2}\mu_{m}^{2}\frac{3p_{0}}{b_{0}}+\widehat{\omega}O(n^{3/2})+in\frac{a_{0}}{b_{0}^{1/2}}+O(n^{1/2})=k_{0}+O(n^{3/2}),\\ k^{\prime}(r_{0})&=(-1+i)O(n^{3/2}),\\ k^{\prime\prime}(r_{0})/2&=k_{2}+n^{3}\mu_{m}\frac{3p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+n^{3}{\widehat{\omega}}\frac{3p_{0}\Lambda^{\prime\prime}(r_{0})}{b_{0}}+O(n^{2}),\end{split}$ (A.4)
@@
Here we verify that, in the case of the trailing vortex (1.3), if $q$ satisfies (1.16), then
@@
Note that this equation has no solution $r>0$ if $\beta\leq q/2$ (in that case (A.8), with “==” replaced by $\geq$”, holds for all $r\geq 0$, with equality only for $r=0$). On the other hand for each $\beta>q/2$ (A.8) has a unique solution $r_{0}>0$. Clearly the solution
$r_{0}=r_{0}(\beta)\quad\text{ is an increasing bijection }r_{0}\colon(q/2,\infty)\to(0,\infty).$ (A.9)
On the other hand, we have $b(r)=4\beta(1-\beta q)q(1-\mathrm{e}^{-r^{2}})\mathrm{e}^{-r^{2}}/(1+\beta^{2}r^{2})$, and $b^{\prime}(r)=0$ if and only if
$g(r)-\beta^{2}=0,\quad\text{ where }g(r):=\frac{1-2\mathrm{e}^{-r^{2}}}{r^{2}(2\mathrm{e}^{-r^{2}}-1)+\mathrm{e}^{-r^{2}}-1}.$ (A.10)
In particular, $b$ has a sign, depending on the sign of $1-\beta q$, and so the third condition in (1.14) implies that we must have $\beta<1/q$. This together with the lower bound on $\beta$ above implies the constraint (1.15).
This in particular gives the first smallness requirement on $q$, namely, $q<2^{1/2}$.
We now observe that $g(r)>0$ if and only if $r<(\log 2)^{1/2}$, and that $g$ is decreasing for such $r$, with $g(r)\to+\infty$ as $r\to 0^{+}$. In order to restrict ourselves to $r\in(0,(\log 2)^{1/2})$, so that we can find such $r$ solving (A.10), we consider $q\in(0,\sqrt{2})$ satisfying (1.16), i.e.
$q<\frac{\log 2}{\sqrt{1-\log 2}}\approx 1.251.$
The last inequality is equivalent to $\beta_{0}<1/q$ where $\beta_{0}>q/2$ is such that $r_{0}(\beta_{0})=\sqrt{\log 2}$, i.e.
$1=\log 2+\frac{\beta_{0}}{q}(\log 2)^{2}$
(note that $\beta_{0}>q/2$ by (A.9)). Thus $g\circ r_{0}\colon(q/2,\beta_{0})\to(0,\infty)$ is a decreasing bijection, and so there exists a unique $\beta\in(q/2,\beta_{0})$ such that
$g(r_{0}(\beta))=\beta^{2}.$
Moreover, $\beta$ is unique also on $(q/2,1/q)$, as $g(r_{0}(\beta))\leq 0$ for $\beta\geq\beta_{0}$. It is also clear that, for such $r_{0},\beta$, we have that $b(r_{0})$ is the global maximum of $b$ and $\Lambda(r_{0})$ is the global minimum of $\Lambda$. Thus we obtain (A.7), as required.
We note that the restriction (1.16) is almost sharp in the sense that a unique choice of $r_{0},\beta$ still exists for some $q>\log 2/\sqrt{1-\log 2}$, but not for $q\to 2^{1/2}$. Indeed, considering $q>1.25$ we also have $\beta^{2}>(q/2)^{2}>0.390625$ and so we could restrict ourselves to the ‘$r$’s for which $g(r)>0.390625$ (since we look for solutions to $g(r)=\beta^{2}$), i.e. to $r\in(0,c):=g^{-1}((0.390625,\infty))\subset g^{-1}((0,\infty))=(0,(\log 2)^{1/2})$. Thus, redefining $\beta_{0}$ to be such that $r_{0}(\beta_{0})=c$, we will have that $\beta<1/q$ if and only if $q$ is less than a number bigger than $\log 2/\sqrt{1-\log 2}$. However, we point out that it is impossible to extend the range of $q$ to the entire interval $(0,\sqrt{2})$, since
$\sup r_{0}((q/2,1/q))\to 0^{+}\quad\text{ as }q\to\sqrt{2}^{-}$
and consequently
$\inf(g\circ r_{0})((q/2,1/q))\to\infty\quad\text{ as }q\to\sqrt{2}^{-}.$
On the other hand, $\beta^{2}\leq 1/q^{2}\leq 1/2$, so, for sufficiently small $|2^{1/2}-q|$, there is no $r_{0}(\beta)$ for which $g\circ r_{0}(\beta)=\beta^{2}$. In other words, there is no solution $r,\beta>0$ to both (A.8) (i.e. $r=r_{0}(\beta)$) and (A.10).